Science Report – April 28th

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: No EVAs took place today.

2. Sunspot and Solar Flare Monitoring: Our CGA continued his task of monitoring the sun for sunspot activity. He found a solar prominence during his imaging session this morning.

3. Shortwave Texting on EVA’s: No EVAs took place today.

4. GPS Route Measurements: No EVAs took place today.

5. Waterless dish cleaning: Samples were taken from a dirty dish that had been place under UV and a dirty dish that had been wiped with dirt and then placed under a UV light. Bacteria are growing on the samples from the first test.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today. Possible that the continuation of this study from Crew 192 will be scrapped.

7. Spacesuit Helmet Fogging: No EVAs took place today.

8. Crew Comfort, Health and Safety on EVA: No EVAs took place today.